OpenStudy (anonymous):

5y = x and 2x -3y = 7 solve using elimination

OpenStudy (anonymous):

2x+5y=7 3x-2y=1 To make the y coefficients subtract out, multiply both sides of the first equation times 2, and the second equation times 5: 2(2x+5y=7) 5(3x-2y=1) 4x+10y=14 15x-10y=5 Add these equations together: 19x=19 x=1

OpenStudy (anonymous):

Substitute x=1 into the first equation: 2x+5y=7 2*1+5y=7 2+5y=7 5y=5 y=1 Substitute both values back into the second equation to check: 3x-2y=1 3*1-2*1=1